OpenStudy (anonymous):

Can someone please help me thank you! solve the equation cosx=-1/2 if 0 ≤ x ≤ 360

OpenStudy (anonymous):

Any help is appreciated thank you

OpenStudy (anonymous):

I believe the answer is 120 degrees and 240 degrees, is that correct?

OpenStudy (anonymous):

cosx = -1/2 x = cos-1 (1/2) = 60 cos is only negative in the second and third quadrant therefore, ITS EITHER x =(180 -60) =120 OR X = (180 + 60) =240

OpenStudy (anonymous):

DEGREES

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

You explained it so well!